### v1

# Backdoor Attack on Vertical Federated Graph Neural Network Learning

Abstract
Federated Graph Neural Network (FedGNN) is a privacy-preserving machine learning technology that combines federated learning (FL) and graph neural networks (GNNs). It offers a privacy-preserving solution for training GNNs using isolated graph data. Vertical Federated Graph Neural Network (VFGNN) is an important branch of FedGNN, where data features and labels are distributed among participants, and each participant has the same sample space. Due to the difficulty of accessing and modifying distributed data and labels, the vulnerability of VFGNN to backdoor attacks remains largely unexplored. In this context, we propose BVG, the first method for backdoor attacks in VFGNN. Without accessing or modifying labels, BVG uses multi-hop triggers and requires only four target class nodes for an effective backdoor attack. Experiments show that BVG achieves high attack success rates (ASR) across three datasets and three different GNN models, with minimal impact on main task accuracy (MTA). We also evaluate several defense methods, further validating the robustness and effectiveness of BVG. This finding also highlights the need for advanced defense mechanisms to counter sophisticated backdoor attacks in practical VFGNN applications.
Keywords FedGNN  $\cdot$ Backdoor Attack  $\cdot$ Graph Neural Network
1 Introduction
Graph Neural Networks (GNNs) are effective tools for handling graph-structured data and have been widely applied to complex problems such as node classification, graph classification, and edge prediction. They have shown great potential in bioinformatics, chemistry, medicine, cybersecurity, linguistics, and transportation [1]. For example, financial risk control systems can train GNNs on transaction data to identify potential fraudulent activities, such as money laundering [2]. However, graph data are often distributed among multiple owners, making direct access impractical due to privacy concerns, regulations, and competitive dynamics [3].
Federated Learning (FL), as a distributed machine learning method that does not require data sharing, effectively solves the data isolation problem. Training GNNs with FL is known as Federated Graph Neural Networks (FedGNN), which can be categorized into two types: Horizontal Federated Learning (HFGNN), where different participants have the same features but different samples [4, 5]; and Vertical Federated Learning (VFGNN), where different participants have the same samples but different features [6, 7]. This paper focuses on VFGNN, which is more common in heterogeneous data environments such as finance and healthcare. Figure 1 illustrates a FedGNN (VFGNN) collaboration for credit analysis involving a fintech company, an online shopping company, and a social media company. The fintech company holds transaction data and labels, the social media company holds social information, and the online shopping company holds shopping information.
Figure 1: A example of VFGNN backdoor attack
Recently, the security implications of FedGNNs have been brought to the forefront of the community. These popular FedGNN have been proven vulnerable to backdoor attacks [8, 9]. Specifically, backdoor attacks against FedGNN implant hidden backdoors into the victim GNN during federated training, which mislead the model on inputs embedded with trigger patterns while having minimal impact on the clean data during inference. In the scenario shown in Figure 1, the social media company can implant a local backdoor to approve loans for users with poor credit ratings, presenting a significant security threat. There has been active research into investigating backdoor attacks tailored to VFL [10, 11, 12], GNN [13, 14, 15, 16], and HFGNN [9, 8]. However, they are not applicable to VFGNN due to unique trigger pattern restrictions derived from both VFL and GNN. To the best of our knowledge, Backdoor attacks on VFGNN are still unexplored, and the robustness of VFGNN under backdoor attacks is essentially unknown.
To fill this gap, this paper proposes a backdoor attack algorithm for Vertical Federated Graph Neural Network Learning, called BVG. The BVG method generates multi-hop triggers based on the graph structure during VFGNN model training, embedding key features of the target class into the model without accessing or modifying labels. Specifically, our contributions in this paper are summarized as follows.
  • To our best knowledge, we propose the first backdoor attack to reveal the vulnerability of VFGNN.
  • We design multi-hop triggers for the graph structure, which inject triggers into the VFGNN without modifying the graph structure.
  • We design a bi-level optimization trigger generation method, which generates triggers based on the graph structure during VFGNN model training. This process creates a trigger containing key features of the target class and integrates it into the model during VFGNN training.
  • Through extensive experiments on widely used datasets, we demonstrate the effectiveness of our proposed method. The BVG method requires knowledge of only 4 target class nodes to achieve a high success rate of backdoor attacks.
2 Related Work
In the fields of VFL and GNN, researchers are investigating backdoor attacks specifically tailored to each domain.
2.1 Backdoor attacks on Vertical Federated Learning
Federated Learning is especially vulnerable to backdoor attacks due to its distributed nature [17]. However, the vertically split model of VFL restricts the attacker’s access to the training labels, let alone modify them.
There are two notable directions in attempting to solve this issue. The first is to conduct label inference attacks [18] to get some labels in advance. For example, VILLAIN [19] leverages label inference to pinpoint samples of the target label and then poisons these samples to inject the backdoor. BadVFL [10] and LR-BA [20] assume attackers can acquire labels of a certain number of samples from each category with label inference. Another direction is to make the knowledge of a small number of labeled samples from the target class a prerequisite [21, 11, 12]. These methods try to use as few as possible (e.g., 500 for [21] and 0.1% for [11]) the labeled target-class samples to establish links between backdoor triggers and target label. Normally, in VFL, the backdoor attack can only be conducted in a clean-label manner [22]. Therefore, works like TPGD [23] assuming the label modification capability of the active party are impractical.
2.2 Backdoor attacks on Graph Neural Networks
The unique challenge to graph-oriented backdoor attacks lies in the inherently unstructured and discrete nature of graph data [14]. For structured and continuous data like images, attackers can directly stamp a trigger pattern (e.g., a black square on the bottom right of the image) $s$ onto a benign image $x$ to achieve a poisoned sample $x_{p}=x+s$ [13]. However, the backdoor attacks against GNN involve design triggers within a large spectrum, including topological structures and descriptive (nodes and edges) features [14]. The effectiveness of backdoor attacks largely depends on designing triggers tailored to the specific attributes of the target task. These triggers mainly fall into three categories: malicious subgraph structures, graph structure perturbations, and node attribute manipulation.
Malicious subgraph structures involve adding malicious subgraphs to nodes [24], edge endpoints [16], or specific positions in the graph [25, 14], causing nodes, edges, or graphs to be misclassified. Graph structure perturbation involves injecting special node connections into the training set adversarially, allowing backdoor attacks to be executed by merely modifying the graph’s topology during the attack [26]. Node attribute manipulation involves carefully designing special node features and adding them to the target class nodes in the training set, followed by adaptive optimization of the graph structure to train a GNN model with backdoors [15, 27]. Although these methods are viable in centralized GNN scenarios, they all rely on full access to the training set, which is nearly impossible in VFGNN.
The aforementioned backdoor techniques, whether tailored for GNN or VFL, are not applicable to VFGNN due to unique threat models. Existing methods such as CBA, DBA, and Bkd-FedGNN focus on HFGNN and do not address the needs of VFGNN [9, 8]. VFL backdoor attack methods lack consideration for the data structure of graph data, while GNN backdoor attacks face difficulties in handling vertically partitioned data. To address these issues, we propose the BVG algorithm for VFGNN, which uses dual-layer optimization and multi-hop triggers to achieve high attack success rates with minimal impact on the main task accuracy (MTA).
3 Proposed Approach
In this section, we formalize the backdoor attack problem in VFGNN and provide a detailed threat model. Next, we present our attack method, with a sketch of the entire approach shown in Figure 2.
3.1 Problem Formulation
Graph data is represented as $G=(V,E,X)$, where $V=\{v_{1},\ldots,v_{n}\}$ is a set of N nodes, $E\subseteq V\times V$ is a set of edges, and $X=\{x_{1},\ldots,x_{N}\}$ is a set of node attributes, with $x_{i}$ being the attribute of node $v_{i}$. The adjacency matrix of graph G is denoted as $A\in\mathbb{R}^{N\times N}$, where $A_{ij}=1$ if nodes $v_{i}$ and $v_{j}$ are connected; otherwise, $A_{ij}=0$.
We focus on the node classification problem, which is common in real-world applications like social networks. In inductive node classification tasks, only a subset of nodes $V_{L}$ in the training graph have labels $Y_{L}=\{y_{1},\ldots,y_{N_{L}}\}$. The test nodes $V_{T}$ are disjoint from the training nodes.
In VFGNN, $K$ ($K>=2$) participants collaboratively train a model using their private data. Each participant $k$ has its local graph $G^{k}=(V,E^{k},X^{k})$, which comprises the entire node set $V$ a subset of edges $E^{k}$, and a subset of feature data $X^{k}$ [6]. From a global perspective, all participants collectively own a global graph $G^{gl}=(V,E^{gl},X^{gl})$, where $E^{gl}=E^{1}\cup\ldots\cup E^{K}$, and for any node attribute in $X$, $x_{k}^{gl}=x_{k}^{1}||\ldots||x_{k}^{K}$, where ||||| | is the concatenation operator. One of the $K$ participants is an active party that knows the corresponding labels of the labeled node set $V_{L}\subseteq V$, and all other participants are negative parties that have no access to the label information.
Each participant employs a local GNN model $f_{k}$ parameterized by $\theta_{k}$ to compute the local output $H^{k}_{l}=f_{k}(\mathcal{G}_{l}^{k};\theta_{k})$, where $\mathcal{G}_{l}^{k}$ denotes the computation graph of node $v_{l}$. In addition to training its own local model, the active party also trains a top model $G$ parameterized by $\theta_{top}$, which aggregates the outputs from all local models to minimize the loss function $\mathcal{L}$. Therefore, the VFGNN model training can be formulated as
$\arg\min_{\Theta}\sum_{v_{l}\in V_{L}}\mathcal{L}(G(H^{1}_{l},\cdots,H^{K}_{l}),y_{l})$ (1)
where $\Theta=\{\theta_{1},\cdots,\theta_{K};\theta_{top}\}$ represents the parameters of the overall VFGNN model.
As an adversary in a backdoor attack, the objective is to establish a mapping association between a trigger and the backdoor target class. This is achieved by injecting a backdoor during the VFGNN training process. Subsequently, upon deployment of the VFGNN model, the adversary can introduce the local trigger to deliberately misclassify the corresponding sample as the target class. The backdoor attack objective in VFGNN is
$\min_{\Theta}\underbrace{\sum_{v_{l}\in V_{L}}\mathcal{L}\left(F(\mathcal{G}_{l};\Theta),y_{l}\right)}_{\text{Main Task}}+\underbrace{\sum_{v_{i}\in V}\mathcal{L}\left(F(a(\mathcal{G}_{i},\delta);\Theta),\tau\right)}_{\text{Backdoor Task}}$ (2)
where $F$ refers to the VFGNN model, which encompasses both the top model and the bottom models of all participating parties. $\delta$ represents the backdoor trigger, a(·) denotes the operation of trigger attachment, and $\tau$ is the backdoor target class.
3.2 Threat model
We assume the adversary is a passive party, not the active party, because the active party can modify labels and thus easily perform backdoor attacks. All other participants are considered trustworthy.
Adversary’s capacity. The adversary adheres to the VFL protocol, transmitting local features and receiving gradients without manipulating other participants’ information.
Adversary’s objective. In VFGNN multi-classification tasks, the adversary’s objective is to inject a backdoor into the model during the training phase. When the poisoned model is deployed for applications, it will misclassify any local data with the backdoor trigger as the designated target class but maintains classification capacity for all other clean data.
Adversary’s knowledge. In backdoor attacks, the adversary requires only a very few training samples labeled target class. Although this requirement deviates from the original VFGNN setting, it is feasible in practical VFGNN applications because it is possible for an adversary to acquire a small number of target class samples through various means, such as direct purchasing [18, 20]. Our experiment results indicate that acquiring only one to four target class nodes is sufficient for a successful backdoor attack. Apart from these limited target samples, the adversary has no information about the models and data of other parties.
Figure 2: A sketch of our proposed Backdoor Attack for VFGNN.
3.3 Multi-hop Trigger Generation
Due to the special nature of the VFGNN architecture, all participants are aware of all nodes in the dataset, but possess different attributes of the nodes. Therefore, the adversary cannot generate a subgraph structure as a trigger like common GNN backdoor attacks [24, 28]. Here, the trigger is added to the node attributes. To prepare a trigger, we propose a multi-hop trigger generation method without introducing new nodes. The trigger is generated and trained with the VFGNN model using bi-level optimization. We use the PGD method [29, 30] to generate the trigger. The optimization formula for the trigger is as follows:
${\delta}_{t+1}=\Pi_{\epsilon}\left(\delta_{t}-\alpha\cdot\operatorname{sgn}\left(\nabla_{\delta}\mathcal{L}\left(F\left(a(\mathcal{G}_{p},{\delta}_{t});\Theta^{*}\right),\tau\right)\right)\right),$ (3)
where $t$ is the step index, $\mathcal{G}_{p}$ represents the computation graph of the node $v_{p}$ to inject the trigger, where $v_{p}\in V_{p}$. $\nabla_{\delta}\mathcal{L}\left(F\left(a(\mathcal{G}_{p},{\delta}_{t});\Theta^{*}\right),\tau\right)$ denotes the gradient of the loss function for the backdoor target class with respect to the trigger, $\alpha$ is the step size, $\Pi_{\epsilon}$ keeps $\delta$ within an $\epsilon$-ball at each step, $F(\Theta^{*})$ refers to the pre-trained VFGNN model. $\operatorname{sgn}(\cdot)$ denotes the sign function.
The multi-hop trigger $\delta$ affects the target node and its neighbors, $\delta=\{\delta^{0},\delta^{1},\cdots,\delta^{M}\}$. Here, $\delta^{m}$ is the trigger added to the attributes of the $m$-hop neighbors of the target node $v_{p}$. Therefore, the operation of adding triggers to $\mathcal{G}_{p}$ can be expressed as
$a(\mathcal{G}_{p},\delta)=(\mathbf{x}_{p}+\delta^{0},\mathbf{X}_{1-hop}+\delta^{1},\cdots,\mathbf{X}_{M-hop}+\delta^{M}),$ (4)
where $\mathbf{x}_{p}$ is the attribute of node $v_{p}$, $\mathbf{X}_{m-hop}$ are the attributes of the $m$-hop neighbors, and $\delta^{0}$, $\delta^{m}$ are the triggers added to these attributes, respectively.
According to (3), the generation of the trigger relies on the gradients of the target class nodes $\tau$. The generation of the trigger and the training of the VFGNN model are accomplished together, which can be expressed through the following bi-level optimization:
$\min_{\delta}\sum_{v_{i}\in\mathcal{V}_{P}}$ $\mathcal{L}(F(a(\mathcal{G}_{i},\delta);\Theta^{*}),\tau)$ (5)
[…]
$+$ $\sum_{v_{i}\in\mathcal{V}_{P}}\mathcal{L}(F(a(G_{i},\delta);\Theta),\tau),$
where the trigger $\delta$ is trained according to all known target class nodes $\mathcal{V}_{P}$ available to the adversary, $\mathcal{V}_{P}\in\mathcal{V}_{L}$. Essentially, this trigger serves as a universal trigger for the target class $\tau$ [31, 32].
Algorithm 1 Multi-hop Trigger Generation
0:  Training dataset $\mathcal{G}$; number of MTG epochs $\mathcal{T}$; target class nodes $\mathcal{V}_{P}$ with target label $\tau$
0:   trigger $\delta$, model parameters $\Theta$
1:  Initialize: $\delta\leftarrow 0$.
2:  while not reached $\mathcal{T}$ do
3:     for $v_{i}$ in $\mathcal{V}_{L}$ parallel do
4:        Adversary $A$: updates $\{\mathcal{G}_{i}^{A}=a(\mathcal{G}_{i}^{A},{\delta})\}_{v_{i}\in\mathcal{V}_{P}}$.
5:        Passive party:
6:        for each party $k=1,2,\dots,K$ parallel do
7:           $k$ computes embedded features $H_{i}^{k}$ using its bottom model $f_{k}$.
8:        end for
9:        Active party:
10:        computes Eq. (1), then updates $\theta_{Top}$ using $\frac{\partial\mathcal{L}}{\partial\theta_{Top}}$.
11:        sends $\frac{\partial\mathcal{L}}{\partial{H}_{i}}$ to all parties.
12:        Adversary $A$:
13:        computes $\nabla_{\delta}\mathcal{L}$ with $\{\frac{\partial\mathcal{L}}{\partial{H}_{i}}\frac{\partial{H}_{i}}{\partial\delta}\}_{v_{i}\in\mathcal{V}_{P}}$
14:        updates $\delta$ with Eq. (3)
15:        Passive party:
16:        for each party $k=1,2,\dots,K$ parallel do
17:           $k$ computes $\nabla_{\theta_{k}}\mathcal{L}=\frac{\partial\mathcal{L}}{\partial H_{i}}\frac{\partial H_{i}^{k}}{\partial\theta_{k}}$.
18:           $k$ updates model parameters $\theta_{k}$.
19:        end for
20:     end for
21:  end while
Algorithm 1 outlines the procedure of the proposed attack. In each iteration of VFGNN model training, the adversary A injects the trigger $\delta$ into the local computation graph of known target class samples $\mathcal{V}_{P}$ according to (4) (line 4). Subsequently, each participant $k$ computes the node embeddings through the bottom model $f_{k}$ (lines 5-8), where $H_{i}^{k}$ represents the embedded features of the $i$-th data from the $k$-th participant. Upon receiving these embedded features, the active party computes the gradients of the loss function with respect to the top model and the embedded features of each participant according to (1). The top model is then updated, and the gradients of the loss with respect to the embedded features $\frac{\partial\mathcal{L}}{\partial{H}_{i}}$ are transmitted to each participant (lines 9-11), where $H_{i}$ denotes the aggregation of the embedded features from all participants for the $i$-th data. After receiving the gradients, the adversary computes $\nabla_{\delta}\mathcal{L}$ and updates the trigger using (3). The current model parameters serve as the pre-trained model $F(\Theta^{*})$ (lines 12-14). Finally, each participant updates its bottom model based on the gradients sent by the active party (lines 15-19).
3.4 Backdoor Trigger and Deployment
After the training is completed, the attacker has successfully injected a backdoor into the model. When the VFGNN model is deployed in practical applications, any local data containing the trigger (added using the method described in (4)) will be misclassified as the specified target class, while other clean data will still be correctly classified. The attacker successfully implements the backdoor attack using a small number of labeled target class samples.
4 Experiments
4.1 Experiments Settings
4.1.1 Local GNN Models
To demonstrate the effectiveness of BVG in VFGNN settings with different local GNN structures, we use three common GNN models as local participants:
Graph Convolutional Network (GCN) [33]: GCN captures local features by propagating and aggregating information between nodes and their neighbors. Each local GNN model in VFGNN uses a two-layer GCN.
Graph Attention Network (GAT) [34]: GAT uses an attention mechanism to assign adaptive weights to neighbor nodes, enhancing the model’s ability to focus on important nodes.
Simple Graph Convolution (SGC) [35]: SGC simplifies GCN by removing nonlinear activations and reducing layers, making it more efficient while retaining essential graph structural information.
To ensure a fair comparison with previous studies [36], we use a two-layer GNN model for each participant’s local GNN to extract local node embeddings, with the dimension set to 16. The number of hidden units is fixed at 32. For GCN and GAT, the activation function is ReLU. VFGNN is trained using Adam, with a learning rate of 0.01.
4.1.2 Datasets
This paper uses three widely adopted public datasets to evaluate the performance of BVG, including Cora [37], Cora_ml [37], and Pubmed [38]. The basic dataset statistics are summarized in Table 1.
Each participant in the VFGNN framework has access to all the nodes in the datasets, but the node features are equally split among the participants. We randomly split the edges of the graphs into equal parts, one for each participant, without overlapped edges between any two participants. We assume the adversary knows only four target class nodes (i.e., $|\mathcal{V}_{p}|=4$), which are randomly selected from the training set.
Datasets Nodes Edges Features Classes
Cora 2708 5429 1433 7
Cora_ml 2810 7981 2879 7
Pubmed 19717 44325 500 3
Table 1: Basic information of the three datasets
4.1.3 Performance Metrics
To evaluate the performance of the BVG method, we employ two common metrics: attack success rate (ASR) and main task accuracy (MTA) [39]. ASR measures the percentage of samples in the backdoor test set that are predicted as the target class by the poisoned model. On the other hand, MTA assesses the accuracy of the clean test set on the poisoned model.
Bottom Datasets MTA ASR
Model Baseline GF TECB BVG Baseline GF TECB BVG
GCN Cora $69.7\pm 2.1$ 48.6$\pm$2.9 62.7$\pm$4.7 65.3$\pm$1.5 8.4$\pm$3.7 51.4$\pm$5.6 16.4$\pm$10.9 $99.9\pm 0.3$
Cora_ml $79.5\pm 2.1$ 64.2$\pm$7.6 77.1$\pm$0.9 76.5$\pm$2.6 4.8$\pm$2.1 26.8$\pm$5.2 34.6$\pm$7.7 $100.0\pm 0.0$
Pubmed $71.7\pm 2.1$ 44.0$\pm$10.2 64.9$\pm$4.2 71.3$\pm$3.5 6.5$\pm$2.2 24.0$\pm$8.0 81.4$\pm$13.8 $100.0\pm 0.0$
GAT Cora $67.8\pm 2.9$ 46.6$\pm$3.9 40.1$\pm$4.1 63.7$\pm$2.2 3.6$\pm$1.2 15.8$\pm$6.3 56.8$\pm$35.5 $99.1\pm 1.1$
Cora_ml $79.5\pm 1.1$ 61.0$\pm$3.3 57.8$\pm$4.9 75.6$\pm$2.8 3.6$\pm$1.6 23.4$\pm$2.7 45.1$\pm$31.0 $99.9\pm 0.2$
Pubmed $74.8\pm 1.0$ 70.8$\pm$3.5 45.0$\pm$17.9 73.3$\pm$1.7 4.5$\pm$0.8 30.2$\pm$2.0 89.6$\pm$20.8 $100.0\pm 0.0$
SGC Cora $68.9\pm 2.3$ 49.0$\pm$3.0 36.7$\pm$30.0 65.3$\pm$2.1 6.6$\pm$5.9 51.0$\pm$4.2 12.2$\pm$13.3 $99.8\pm 0.4$
Cora_ml $79.5\pm 1.4$ 67.4$\pm$4.0 74.8$\pm$2.6 74.9$\pm$2.4 4.6$\pm$2.5 27.8$\pm$4.0 15.2$\pm$9.1 $100.0\pm 0.0$
Pubmed $72.1\pm 2.0$ 41.8$\pm$12.3 54.0$\pm$27.0 70.6$\pm$2.9 7.1$\pm$1.0 32.7$\pm$10.9 59.8$\pm$48.8 $99.9\pm 0.2$
Table 2: Performance comparison with other methods on three datasets and three bottom models (Standard Deviation Included, Best Results Highlighted in Bold.)
4.1.4 Comparison Benchmarks
Due to the limited research on backdoor attacks against VFGNN, we chose two attack methods from other scenarios and the baseline VFGNN (without backdoors) for comparison. To make the attack methods from other scenarios transferable to VFGNN, we modified the comparison methods to a certain extent to match the VFGNN scenario and tasks. The comparison methods are briefly described as follows:
  • GF [36]: This method is designed for adversarial tasks on VFGNN. The original adversarial goal is to maximize the loss of node classification. To match our attack goal, we modified the adversarial goal to minimize the loss of the backdoor class. This method is mainly divided into three parts: obtaining the embedding vectors of other participants through queries, training a proxy model using the obtained embeddings, and obtaining the adversarial graph on the proxy model.
  • TECB [11]: This method is designed for backdoor attacks in VFL, mainly targeting image data. To adapt it to backdoor tasks on graphs, we set the target node’s attributes as the goal of its trigger embedding. This method is mainly divided into two stages: first, using limited known labels to train the trigger, and then using the trigger for target gradient alignment.
4.2 Experiment Results Analysis
4.2.1 Attack Efficacy in Two-Party Setting
We first conducted experiments in a VFGNN framework with two parties, and the obtained experimental results are shown in Table 2. To evaluate the impact of the BVG method on the main classification task, we also tested the performance of VFGNN without any backdoor attack (the Baseline in Table 2). The table shows that both comparison methods and BVG have a decline in the main task performance compared to the baseline, but BVG has the smallest decline, thus achieving the highest MTA among the attack methods. Table 2 also shows that BVG has a significantly higher backdoor attack success rate than those of to the compared methods for all the tested GNN models and used datasets. These results indicate that BVG achieves a high attack success rate with the least impact on the main task performance, making it an effective backdoor attack method to VFGNN.
4.2.2 Attack Efficacy in Multi-Party Setting
We also evaluated the proposed attack method’s performance in multi-party VFGNN scenarios where multiple participants are involved. Unlike the two-party setting, multi-party scenarios become more complex due to the increased number of participants and their interactions. Our evaluation focuses on how the presence of multiple participants affects the attack success rate and the main task performance of VFGNN.
We conducted experiments with two, three, and four participants to evaluate the scalability and effectiveness of the proposed backdoor attack method in multi-party VFGNN scenarios. Each participant was assigned a subset of the graph’s edges and node features, with one active party holding the labels for the labeled node set $V_{L}$. The passive participants, including the adversary, only had partial information about the graph. The experimental results are shown in Figure 3.
The results indicate that the main task accuracy slightly decreases as the number of participants increases. This trend could be attributed to the increased complexity and noise introduced by multiple participants. The backdoor attack success rate remains high across all multi-party settings but slightly decreases as the number of participants increases. The obtained results verify that our attack method is effective even in more complex multi-party environments. The slight decrease in success rate may be due to the dilution of the adversary’s influence with more participants.
Figure 3: Performance of BVG under Multi-Party Settings (Standard Deviation Included.)
Figure 4: Evaluation of BVG under Defenses: The ISO, and GC Methods Demonstrate Protection Performance Against BVG with Various Noise Scales (0.5, 0.1, 0.01, 0.001) and Compression Rates (0.75, 0.5, 0.25, 0.1). DP Noise Scales are (0.75, 0.5, 0.25, 0.1). Max Norm Methods are Depicted as Single Points. The Optimal Position for a Defense Method is the Lower Right Corner, Indicating the Highest MTA and the Lowest ASR.
4.2.3 Attack Efficacy under Defense
To evaluate the robustness of the BVG method, we explored potential backdoor defense strategies within the VFGNN framework. We considered two main types of defense: the first type is GNN backdoor defense methods, such as Prune and Prune+LD [24]. These methods defend against backdoor attacks by pruning the edges between nodes with low similarity. Pruning these edges can disrupt the attacker’s trigger structure and connections. However, VFGNN’s architecture prevents defenders from pruning the adversary’s data, making these methods unsuitable. The second type is VFL backdoor defense methods, which reduce the leakage of labels, features, and other information during training by perturbing the gradients during the VFGNN training process, thereby lowering the likelihood of attackers implanting backdoors.
We implemented several representative VFL backdoor defense methods within the VFGNN framework. These defenses include Differential Privacy (DP) [18], max norm [40], Gradient Compression (GC) [18, 41], and adding Gaussian noise to intermediate features. These methods defend against backdoor attacks by perturbing the gradients sent from the active party to the adversary during VFGNN training. In our experiments, we set the compression rates of GC and the scale of DP to 0.75, 0.5, 0.25, and 0.1, and the scale of Gaussian noise to 0.5, 0.1, 0.01, and 0.001.
The obtained results are depicted in Figure 4 and show a trade-off between model utility and backdoor protection in VFGNN. Although different defense strategies can partially reduce the impact of backdoor tasks, they also degrade the performance of the main task. In our experiments with three underlying models and three datasets, ISO provided a relatively effective defense at a noise scale of 1, reducing the backdoor effect across all three datasets but at the cost of significantly lowering the accuracy of the main task. Therefore, the results indicate that typical defense methods cannot easily mitigate BVG attacks without severely sacrificing the performance of the main task, validating the robustness of the BVG method in resisting VFGNN defenses.
4.3 Analysis of Hyperparameter in BVG
4.3.1 Impact of Multi-hop Trigger in BVG
Bottom Datasets ASR ASR ASR
Model (0 hop) (1 hop) (2 hops)
GCN Cora 74.13 98.23 98.83
Cora_ml 58.30 99.80 99.70
Pubmed 80.13 100.00 100.00
GAT Cora 61.27 89.47 100.00
Cora_ml 41.23 86.00 99.47
Pubmed 89.08 100.00 100.00
SGC Cora 89.08 99.14 99.76
Cora_ml 59.53 97.20 99.67
Pubmed 93.32 99.50 99.90
Table 3: Performance Impact of Multi-Hop Triggers on BVG Method (Best results are highlighted in bold.)
Considering the sensitivity of GNN models to spatial relationships, we employed multi-hop triggers in the BVG method. This involves injecting triggers into the multi-hop neighbors of the target node to enhance the performance of backdoor attacks. Since the Bottom model used in our experiment is a 2-layer GNN, its receptive field is limited to 2-hop neighbors. Therefore, we only considered the trigger injection up to 2 hops. To evaluate the impact of multi-hop triggers on BVG performance, we compared the backdoor task performance of the BVG method under three scenarios: without using multi-hop triggers (0 hop), considering only 1-hop neighbors (1 hop), and considering 2-hop neighbors (2 hops). The results are shown in Table 3.
We also conducted experiments on three datasets using three bottom models, with each value being the average of five independent trials. It can be observed that the more hops considered for the trigger, the better the ASR performance. Compared to not using multi-hop triggers, the performance improvement when using 2-hop triggers can even exceed 20%. This demonstrates the effectiveness of multi-hop triggers for VFGNN backdoor attacks.
4.3.2 Impact of Target Dataset Size
The BVG method requires a minimal amount of auxiliary target sample labels for backdoor attacks. These labeled nodes play a crucial role in BVG, making the size of $\mathcal{V}_{p}$ (denoted as $N$) an important hyperparameter that affects BVG’s effectiveness and its real-world threat. We measured the impact of different sizes of $\mathcal{V}_{p}$ on BVG performance (ASR and MTA) across three base models and three datasets, with the results plotted in Figure 5.
Figure 5: Impact of Different Sizes of $\mathcal{V}_{p}$ on BVG Performance (ASR and MTA) Across Three Datasets.
The figure shows that the BVG method can achieve excellent backdoor attack performance even with very limited auxiliary target data. In the Cora_ml and Pubmed datasets, the BVG method performs well with just one target class node. For more complex datasets like Cora, the method achieves over 80% attack accuracy with just one target class node, and only four target class nodes are needed to achieve good ASR and MTA performance. This demonstrates that BVG can effectively carry out backdoor attacks in VFGNN, even with very limited information about target samples, thus providing an efficient attack method in practical VFGNN scenarios.
4.3.3 Impact of Trigger Perturbation
In backdoor attacks, trigger perturbations are usually constrained to ensure stealthiness. Figure 6 shows the performance of BVG on three datasets and three Bottom models with perturbations ($\epsilon$) of 0.002, 0.02, and 0.2.
Figure 6: Performance of BVG on Three Datasets and Three Bottom Models with Different Perturbation Levels ($\epsilon$)
We conducted experiments on three datasets and observed that as the perturbation increases, the ASR performance of the backdoor attack significantly improves. The required perturbation value varies across different datasets, with the Cora dataset requiring a larger perturbation than the other datasets. Therefore, within the constraints of ensuring stealthiness, selecting an appropriate range of perturbation values can achieve a higher success rate for backdoor attacks.
5 Conclusion
In this paper, we proposed the BVG method for backdoor attacks to Vertical Federated Graph Neural Network Learning (VFGNN). Utilizing a multi-hop trigger generation approach, the BVG method can perform efficient backdoor attacks with very limited knowledge of target class nodes. Extensive experiments using three GNN models across three datasets indicate that BVG achieves high attack success rates (ASR) with minimal impact on the main task accuracy (MTA). The evaluation of BVG efficacy under various defense methods highlights the robustness and efficiency of the proposed attack method, underscoring the necessity for advanced defense mechanisms in practical federated learning applications to counter such sophisticated backdoor attacks.